OpenStudy (anonymous):

Calc 3: Find c such that, a=ci- 2j and b= -4i+3j are parallel. (i and j are unit vectors). Can someone help me on this too? -__-

OpenStudy (kinggeorge):

To find parallel vectors, it's pretty straightforward. All you need to know is that if you have two parallel vectors \[\vec{r_1}=a\,\vec{i}+b\,\vec{j}\]\[\vec{r_2}=c\,\vec{i}+d\,\vec{j}\]then\[\frac{a}{c}=\frac{b}{d}\]

OpenStudy (kinggeorge):

This means you can set up the proportion \[\frac{-2}{3}=\frac{c}{-4}\]and solve for \(c\).

OpenStudy (anonymous):

thank you

OpenStudy (kinggeorge):

You're welcome.